Each of 8

 students reported the number of movies they saw in the past year. This is what they reported:8, 5, 16, 12, 17, 9, 5, 13

Find the mean and median number of movies that the students saw.

If necessary, round your answers to the nearest t


1
Expert's answer
2021-03-16T10:19:17-0400
"5,5,8,9,12,13,16,17"

"mean=\dfrac{5+5+8+9+12+13+16+17}{8}"

"=10.625\approx10.6"


"median=\dfrac{9+12}{2}=10.5"